# An Explicit Wavefunction of the Interacting Non-Hermitian Spin-1/2 1D System

### v1

@@ Abstract
We present an explicit Bethe-ansatz wavefunction to a 1D spin-$\frac{1}{2}$ interacting fermion system, manifesting a many-body resonance resulting from the interplay between interaction and non-Hermitian spin-orbit coupling. In the dilute limit, the wavefunction is greatly simplified and then factorized into Slater determinants and a Jastrow factor. An effective thermodynamic distribution is constructed with an effective Hamiltonian including a repulsion resulting from Pauli’s exclusion principle and a distinctive zigzag potential arising from the resonance. The competition between these effects leads to a transition from a uniformly distributed configuration to a phase separation. The connection to the recent cold atom experimental efforts of realizing on-site atom-loss is discussed.
@@
Introduction.— The non-Hermitian skin effect (NHSE) has attracted significant attentions in recent years [1, 2, 3, 4, 5]. The associated anomalous properties, such as the complex-valued spectrum and the localization on boundaries, can be described by the theory of the generalized Brillouin zone, in which momenta are complex-valued [2, 6, 7, 8, 9, 10, 11]. These distinctive features are highly sensitive to boundary conditions. For example, the eigenstates in non-Hermitian systems are usually extended under the periodical boundary condition (PBC) while localized under the open boundary condition (OBC), which contrasts to the case in Hermitian physics. Experimentally, the NHSE has been observed in various systems, including metamaterials [12, 13], photonic systems [14], electrical circuits [15, 16], acoustic crystals [17], and cold atomic systems [18].
@@
Despite the significant progress in the NHSE, current studies predominantly focus on the single-body physics. How the NHSE behaves under strong interactions remains an open question, and non-perturbative analytical studies are desired. The Bethe-ansatz (BA) method [19, 20] is a systematic tool for studying one dimensional (1D) integrable systems, including the Lieb-Liniger model of the interacting Bose gas [21, 22], the Gaudin-Yang model of the interacting Fermi gas [23, 24], and the Lieb-Wu solution to the Hubbard model [25]. When applied to non-Hermitian systems, it has been found that NHSE is suppressed by repulsive interactions [26, 27, 28, 29, 30]. However, the complexity of BA wavefunctions makes it difficult to calculate observables. It would be desired to construct an explicit many-body wavefunction to facilitate a deeper understanding of the NHSE in interacting systems, akin to the Ogata-Shiba-type and the Laughlin-type wavefunctions [31, 32].
@@
In this work, we present a concise expression for the many-body wavefunction in a 1D spin-1/2 fermion system with the non-Hermitian spin-orbit coupling (SOC). As a result of the strong repulsive $\delta$-interaction, each particle behaves as a soft boundary to particles with opposite spins, which induces an effective attraction between them. Resonance states are formed here instead of bound states, i.e., the Bethe string states [33, 34, 35]. The explicit many-body wavefunction is constructed in the dilute limit, which is a rare example in many-body physics. It consists of the product of Slater determinants and the Jastrow factor [36] reflecting the resonance between particles with opposite spins. Remarkably, as varying the interacting strength, a phase transition takes place that an effective “spin-diople” per resonance pair scales from a finite value to a linear divergence with the system size.
@@
$\hat{H}=\sum_{l=1}^{N}\frac{\left(-i\nabla_{l}+im\alpha\sigma_{l}^{z}\right)^{2}}{2m}+\sum_{\langle ll^{\prime}\rangle}2g\delta\left(x_{l}-x_{l^{\prime}}\right),$ (1)
where $\alpha>0$ represents the strength of the non-Hermitian SOC; $g>0$ represents the strength of repulsive interaction; the PBC is assumed. Since the $z$-component of total spin is conserved, the eigenvalues and many-body eigenstates can be labelled by the particle numbers of two components, i.e., $N_{\uparrow}$ and $N_{\downarrow}$. In the following, the real and imaginary parts of the complex momentum are defined as
@@
where $l$ is the particle number index and $\sigma_{l}$ is the spin component along the $z$-direction.
We warm up by considering the single-body problem. If the particle carries spin $\sigma$ ($\sigma=\uparrow,\downarrow$), the eigenstates are denoted as $e^{ik_{\sigma}x}|\sigma\rangle$, and the eigen-energies are $(k_{\uparrow,\downarrow}\pm im\alpha)^{2}/(2m)$, respectively. The momentum is quantized as $k_{\sigma}=2\pi n_{\sigma}/L$ under the PBC. Upon the OBC, the spin-up and down particles localize at the right and left boundaries, respectively. The single-particle localization length is $\lambda_{s}=(m\alpha)^{-1}$, which is independent of the system size, indicating the presence of bound states. This is the conventional NHSE discussed in the literature.
Figure 1: The 2-body wavefunction $\varphi(x_{1},x_{2})$ with particles of opposite spins by fixing the coordinate of the spin-up one $x_{2}=0$. The parameter values are $\chi_{1}=\chi_{2}=0$, $m=1$, $g=2$, $\alpha=1$ and $L=50$. The peak appears at $x_{1}\sim\ln(L\lambda_{s}^{-1})/(L\lambda_{s}^{-1})$. In the case of $L\gg\lambda_{s}$, the peak is located at $x_{1}=x_{2}$ and $\varphi$ becomes discontinuous. It implies that the spin-down particle tends to lie on the right side of the spin-up particle, forming a resonant pair on the ring.
@@
$\left\{\begin{array}[]{l}\left(-\frac{\nabla_{X}^{2}}{4m}-m\alpha^{2}\right)\Phi(X)=E_{X}\Phi(X),\vspace{1ex}\\ \left(-\frac{\nabla_{r}^{2}}{m}-2\alpha\nabla_{r}+2g\delta(r)\right)\phi(r)=E_{r}\phi(r).\end{array}\right.$ (3)
$\Phi(X)$ is solved as $e^{iK_{X}X}$ where $K_{X}=2\pi n_{K}/L$. The non-Hermitian term $-2\alpha\nabla_{r}$ only appears in the motion of the relative coordinate, where the $\delta$-potential acts as a soft boundary. Therefore, the reminiscence of the NHSE would bring an effective attraction between two particles, explained as follows. The relative motion is solved as
@@
where $\beta=\alpha+i\frac{k_{r}}{m}$. As shown in Supplemental Material (SM) A, in the case of $L\gg\lambda_{s}$, Eq. (5) is solved as
$\left\{\begin{array}[]{l}\chi_{r}=(2n_{r}+n_{K})\pi,\ \ \ \mathcal{\eta}_{r}=\ln(1+\frac{g}{\alpha}),\vspace{1ex}\\ \frac{A}{B}=-(1+\frac{\alpha}{g}).\end{array}\right.$ (8)
$\phi(r)$ is identical to a single-body wavefunction of spinless particle subjected to a “soft” boundary condition, which lies between the cases of OBC and PBC, since the $\delta$-potential permits partial transmission. In all cases, the solution possesses a pair of momenta, whose imaginary parts are summed to $2m\alpha$. In the OBC case, both imaginary parts equal $m\alpha$, while in the PBC case, one becomes real and the other carries the imaginary part of $2m\alpha$. In our case, a small imaginary part $\eta_{r}/L$ is at the order of $1/L$, and the other remains at the order of $2m\alpha$. Consequently, the decay of $\phi(r)$ is at the length scale of $L$, such that these states are resonance rather than bound states [37, 38]. The situation of OBC is recovered for $g\sim\frac{1}{mL}e^{L/\lambda_{s}}$ in which case the localization length $L/\eta_{r}\sim\lambda_{s}$, while that of PBC corresponds to $g=0$.
@@
where $0<x_{1,2}<L$. $B$ and $B^{\prime}$-terms are the reflected waves of $A$ and $A^{\prime}$-terms respectively. After the collision, the real parts of momenta switch, but their imaginary parts change due to the SOC. For example, in the domain of $\theta(x_{1}>x_{2})$, the imaginary parts of $k_{1,\downarrow}$ and $k_{2,\uparrow}$ are at the order of $1/L$,
@@
where $\chi_{i}=2\pi n_{i}$ with $i=1,2$. In contrast, the imaginary parts of the reflected momenta $k_{1,\uparrow}$ and $k_{2,\downarrow}$ become finite as
@@
$A^{\prime}$ and $B^{\prime}$-terms are the transmitted waves of $A$ and $B$-terms respectively. The PBC yields $A^{\prime}/A=e^{ik_{1,\downarrow}L}$ and $B^{\prime}/B=e^{ik_{2,\downarrow}L}$.
We view $A$ and $A^{\prime}$-terms as the incident waves and $B$ and $B^{\prime}$-terms as the reflection ones. Due to the different behaviors of their imaginary parts, in the case of $L\gg\lambda_{s}$, the reflection channels can be dropped if the inter-particle distance exceeds $\lambda_{s}$. Then the wavefunction is simplified to the product of plane-waves and a Jastrow factor,
@@
where $e^{-\frac{1}{2}W(x_{1}-x_{2})}$ is given by the sum of step-functions modified by the imaginary parts of the complex momentum
@@
$W(r)=\begin{cases}2\eta_{r}\frac{r}{L},&0<r<L\vspace{1ex}\\ 2\eta_{r}(\frac{r}{L}+1),&-L<r<0\end{cases},$ (13)
which exhibits a jump at $r=0$.
The exact and approximated wavefunctions Eq. (9) and Eq. (12) are illustrated in Fig. 1 respectively. The spin-up particle is fixed at $x_{2}=0$. Increasing $x_{1}$ from 00, $\varphi(x_{1},x_{2})$ rapidly reaches the peak located at $x_{1,\rm peak}\sim\ln(L\lambda_{s}^{-1})/(L\lambda_{s}^{-1})$. If $L\gg\lambda_{s}$, $x_{1,\rm peak}$ coincides with $x_{2}$. The peak is followed by a slow decay at the length scale of $L$. The behavior at $x_{1}<0$ can be obtained by applying the PBC. The spin-down particle prefers the right side of the spin-up one, which can be understood as a weaker version of the NHSE.
The energy eigenvalue is $E=(k_{1,\downarrow}-im\alpha)^{2}/2m+(k_{2,\uparrow}+im\alpha)^{2}/2m$, where $k_{1,\downarrow}$and $k_{2,\uparrow}$ are complex. The $\delta$-interaction only modifies the allowed values of the momenta, whose imaginary parts always tend to reduce the effect of non-Hermitian SOC. As we will explain later, the two-body wavefunction can be generalized to the many-body case, in which the above picture still holds.
Many-body problem: BA equations.— For eigenstates with $N_{\downarrow}$ down spins and $N_{\uparrow}=N-N_{\downarrow}$ up spins, the corresponding BA equations are [39]:
@@
It would be challenging to solve $\{\bar{k}_{l}\}$ from the BA equations. Actually it is significantly simplified due to the non-Hermitian SOC. Among all the scattering channels, there is a specific one with momentum distribution
$k_{i,\downarrow}$ $=$ $(\chi_{i}+i\mathcal{\eta}_{r}N_{\uparrow})/L,~{}~{}k_{j,\uparrow}=(\chi_{j}-i\mathcal{\eta}_{r}N_{\downarrow})/L,$ (16)
where $1\leq i\leq N_{\downarrow}$, $N_{\downarrow}+1\leq j\leq N$, $\chi_{l}=2\pi n_{l}$ and $\mathcal{\eta}_{r}$ is defined in Eq. (8). We denote this channel as the incident wave, in which the imaginary parts of momenta are small in the dilute limit. Similar to the two-body case, reflection between spin-up and down particles will cause a pair of momenta to carry finite imaginary parts at the order of $m\alpha=\lambda_{s}^{-1}$. These reflected waves decay much faster than the incident wave, therefore can be discarded in the dilute limit defined as,
@@
where $d=L/N$ is the average inter-particle distance. Detailed calculations are found in the SM. B. Note that our approximated solution exhibits a singularity at $\alpha=0$, since the dilute limit is broken in that case.
@@
Figure 2: ($a$) and ($b$) plots $V(r)$ and $W(r)$, the effective potentials arising from transforming the BA wavefunction into a thermodynamic distribution shown in Eq. (20). Configuration ($c$) or ($d$) is favored when $V$ or $W$ is dominant, respectively. The former exhibits a finite value of dipole strength $p$, while $p$ diverges linearly with $L$ in the latter case. The Monte-Carlo simulations of $p$ at $L=100$ with different particle numbers are presented in ($f$), which shows a transition from ($c$) to ($d$) as increasing $\eta_{r}$ defined in Eq. (8). A jump of $p$ shows that the transition is of 1st order.
@@
$\varphi(x_{\downarrow};x_{\uparrow})$ $=\!\text{det}(e^{i\chi_{i_{1}}x_{i_{2}}/{L}})\text{det}(e^{i\chi_{j_{1}}x_{j_{2}}/{L}})$ (18)
×$\times e^{-\frac{1}{2}\sum_{ij}\!W(x_{i}-x_{j})},$
in which two $\text{det}(\cdots)$ represent the Slater determinants of the plane-waves for spin-up and down particles respectively; the two-body Jastrow factor in Eq. (12) is also generalized to the many-body case. Here $1\leq i,i_{1},i_{2}\leq N_{\downarrow}$ and $N_{\downarrow}+1\leq j,j_{1},j_{2}\leq N$ are the coordinate indexes for spin-down and up particles respectively. This wavefunction is similar to that of the Hubbard model at $U\rightarrow\infty$ [31], where the BA wavefunction is factorized into a Slater determinaint of spinless fermions and a BA wavefunction of the spin-1/2 chain.
The above simplification is justified as follows. For a given coordinate permutation, the wavefunction is organized into different Slater determinants, corresponding to different scattering channels. As explained before, the one denoted as incident wave dominates over its reflection descendants in the dilute limit. As a good approximation, the wavefunction is a summation of these incident waves with different coordinate permutations. In other words,
$\varphi(x_{\downarrow};x_{\uparrow})$ $=\text{det}(e^{i\chi_{i_{1}}x_{i_{2}}/{L}})\text{det}(e^{i\chi_{j_{1}}x_{j_{2}}/L})e^{-\frac{1}{L}\sum_{ij}\eta_{r}(x_{i}-x_{j})}$
×$\times\sum_{Q}\theta\left(x_{Q_{1}}>x_{Q_{2}}>\cdots>x_{Q_{N}}\right)A(Q).$
where $Q$ represents the permutation $x_{Q_{1}}>x_{Q_{2}}>\cdots>x_{Q_{N}}$ with $1\leq Q_{l}\leq N$. $\sum_{Q}$ denotes the sum over all permutations. Different incident waves are connected by transmissions with a “phase shift”, whose module is not $1$ due to the non-Hermitian SOC. After switching a pair of neighboring particles with opposite spins, the amplitudes are changed by
@@
which is momentum independent at the leading order. As proved in SM. B, in this case the summation of step functions can be organized into
@@
$-\frac{N-2}{4},-\frac{N-6}{4},\cdots-1,0,1,\cdots\frac{N-6}{4},\frac{N-2}{4}.$
@@
$\mathcal{H}$ $=$ $\sum_{i_{1}<i_{2}\in\{\downarrow\}}\!V(x_{i_{1}}\!-x_{i_{2}})+\!\!\sum_{j_{1}<j_{2}\in\{\uparrow\}}\!V(x_{j_{1}}\!-x_{j_{2}})$ (21)
$+$ $\sum_{i\in\{\downarrow\},j\in\{\uparrow\}}W(x_{i}-x_{j}),$
in which $V(r)=-2\ln\left|\sin\frac{\pi}{L}r\right|$ originates from the Pauli exclusion principle. $V(r)$ and $W(r)$ are plotted in Fig. 2 ($a$) and ($b$) respectively. $V$ describes an effective repulsion between particles of identical spins, while $W$ brings an unidirectional attraction between opposite spins, with spin-up particles preferring the left side of spin-down ones.
If $V$ dominates, particles tend to uniformly distribute along the ring, with a weak pairing tendency between opposite spins to take the advantage of $W$, as depicted in Fig. 2 ($c$). Conversely, if $W$ dominates, the system prefers phase separation, such that nearly all spin-up particles lie on the left of spin-down particles, making the configuration in Fig. 2 ($d$) more favorable. The competition between $V$ and $W$ is investigated via Monte-Carlo simulations with the probability distribution Eq. (20). We consider the “dipole” strength $p=\frac{1}{(N/2)^{2}}\sum_{ij}p_{ij}$, where $p_{ij}$ is the dipole between a spin-down and up particle located at $x_{i\downarrow}$ and $x_{j\uparrow}$ respectively,
$p_{ij}=\left\{\begin{array}[]{l}r_{ij},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}|r_{ij}|\leq L/2,\\ r_{ij}-\operatorname{sgn}(r_{ij})L,\ |r_{ij}|>L/2,\end{array}\right.$ (22)
with $r_{ij}=x_{i\downarrow}-x_{j\uparrow}$. Within this definition, $p$ ranges from $-L/2$ to $L/2$. The thermodynamics average $\langle p\rangle$ is plotted in Fig. 2 ($f$). As $\eta_{r}$ defined in Eq. (8) increases, $\langle p\rangle$ evolves from a finite value, which is consistent with Fig. 2 ($c$), to a linear divergence with $L$, illustrated in Fig. 2 ($d$). The transition takes places at $\eta_{r}\approx 3.15$.

### v2

@@ Abstract
We present an explicit Bethe-ansatz wavefunction to a 1D spin-$\frac{1}{2}$ interacting fermion system, manifesting a many-body resonance resulting from the interplay between interaction and non-Hermitian spin-orbit coupling. In the dilute limit, the Bethe-ansatz wavefunction is factorized into Slater determinants and a Jastrow factor. An effective thermodynamic distribution is constructed with an effective Hamiltonian including a repulsion resulting from Pauli’s exclusion principle and a distinctive zigzag potential arising from the resonance. The competition between these effects leads to a transition from a uniformly distributed configuration to a phase separation. Clustering of particles with identical spins is observed in the latter phase, demonstrating that the many-body resonance effect is enhanced by the repulsive interaction.
@@
Introduction.— The non-Hermitian skin effect (NHSE) has attracted significant attentions recently [18, 28, 42, 16, 17, 1, 34, 26]. The associated exotic properties, such as the complex-valued spectrum and the localization on boundaries, can be described by the theory of the generalized Brillouin zone, in which momenta are complex-valued [42, 14, 44, 40, 41, 31, 45]. These distinctive features are highly sensitive to boundary conditions. For example, the eigenstates in non-Hermitian systems are extended under the periodical boundary condition (PBC) while localized under the open boundary condition (OBC), which contrasts to the case in Hermitian physics. Experimentally, the NHSE has been observed in various systems, including metamaterials [2, 7], photonic systems [37], electrical circuits [10, 25], acoustic crystals [46], and cold atomic systems [21].
@@
Despite the significant progress in the NHSE, current studies predominantly focus on the single-body physics. How the NHSE behaves under strong interactions remains an open question, and non-perturbative analytical studies are desired. The Bethe-ansatz (BA) method [13, 35] is a systematic tool for studying one dimensional (1D) integrable systems, including the Lieb-Liniger model of the interacting Bose gas [22, 24], the Gaudin-Yang model of the interacting Fermi gas [38, 6], and the Lieb-Wu solution to the Hubbard model [23]. When applied to non-Hermitian systems [48, 4, 5], it has been found that NHSE is suppressed by repulsive interactions [29, 27, 33, 12, 47]. However, the complexity of BA wavefunctions makes it difficult to calculate observables. It would be desired to construct an explicit many-body wavefunction to facilitate a deeper understanding of the NHSE in interacting systems, akin to the Ogata-Shiba-type and the Laughlin-type wavefunctions [30, 15, 8].
@@
In this work, we present a concise expression for the many-body wavefunction in a 1D spin-1/2 fermion system with the non-Hermitian spin-orbit coupling (SOC). As a result of the repulsive $\delta$-interaction, each particle behaves as a soft boundary to particles with opposite spins, inducing an effective unidirectional attraction between them. Resonance states are formed here instead of bound states, i.e., the Bethe string states [39, 36, 32]. The explicit many-body wavefunction is constructed in the dilute limit, which is a rare example in many-body physics. It consists of the product of Slater determinants and the Jastrow factor [11] reflecting the resonance between particles with opposite spins. Remarkably, a phase transition occurs as the interacting strength increases, where the level of localization exhibits a jump, indicating that the resonance is enhanced by the repulsive interaction.
@@
$\hat{H}=\sum_{l=1}^{N}\frac{\left(-i\nabla_{l}+im\alpha\sigma_{l}^{z}\right)^{2}}{2m}+\sum_{l<l^{\prime}}2g\delta\left(x_{l}-x_{l^{\prime}}\right),$ (1)
where $\alpha>0$ and $g>0$ represent the strength of the non-Hermitian SOC and the repulsive interaction respectively; the PBC is assumed. Since the $z$-component of total spin is conserved, the eigenvalues and eigenstates can be labelled by the particle numbers of two components, i.e., $N_{\uparrow}$ and $N_{\downarrow}$. In the following, the real and imaginary parts of the complex momentum are defined as
@@
where $\sigma_{l}=\pm 1$ represents the spin $z$-component.
We warm up by considering the single-body problem. If the particle carries spin $\sigma$ ($\sigma=\uparrow,\downarrow$), the eigenstates are denoted as $e^{ik_{\sigma}x}|\sigma\rangle$, and the eigen-energies are $(k_{\uparrow,\downarrow}\pm im\alpha)^{2}/(2m)$. The momentum is quantized as $k_{\sigma}=2\pi n_{\sigma}/L$ under the PBC. Upon the OBC, the spin-up and down particles localize at the right and left boundaries, respectively. The single-particle localization length is $\lambda_{s}=(m\alpha)^{-1}$, which is independent of the system size, indicating the presence of bound states. This is the conventional NHSE discussed in the literature.
@@
$\begin{array}[]{l}\left(-\frac{\nabla_{X}^{2}}{4m}-m\alpha^{2}\right)\Phi(X)=E_{X}\Phi(X),\vskip 4.30554pt\\ \left(-\frac{\nabla_{r}^{2}}{m}-2\alpha\nabla_{r}+2g\delta(r)\right)\phi(r)=E_{r}\phi(r).\end{array}.$ (3)
$\Phi(X)$ is solved as $e^{iK_{X}X}$ where $K_{X}\!=\!2\pi n_{K}/L$. The non-Hermitian term $2\alpha\nabla_{r}$ only appears in the motion of the relative coordinate, where the $\delta$-potential acts as a soft boundary. As a result, the reminiscence of the NHSE would bring an effective unidirectional attraction between two particles, explained as follows. The relative motion is solved as
@@
where $\beta=\alpha+i\frac{k_{r}}{m}$. As shown in Supplemental Material (SM) I, in the case of $L\gg\lambda_{s}$, Eq. (5) is solved as
$\!\!\!\!\chi_{r}=(2n_{r}+n_{K})\pi,\ \mathcal{\eta}_{r}=\ln(1+\frac{g}{\alpha}),\ \frac{A}{B}=-(1+\frac{\alpha}{g}).$ (6)
$\phi(r)$ is identical to a single-body wavefunction of spinless particle subjected to a ‘soft’ boundary condition, which lies between the cases of OBC and PBC, since the $\delta$-potential permits partial transmission. In all cases, the solution possesses a pair of momenta, whose imaginary parts are summed to $2m\alpha$. In the OBC case, both imaginary parts equal $m\alpha$, while in the PBC case, one becomes real and the other carries the imaginary part of $2m\alpha$. In our case, a small imaginary part $\eta_{r}/L$ is at the order of $1/L$, and the other remains at the order of $2m\alpha$. Consequently, the decay of $\phi(r)$ is at the length scale of $L$, such that these states are resonance rather than bound states [20, 19]. The situation of OBC is recovered for $g\sim\frac{1}{mL}e^{L/\lambda_{s}}$ in which case the localization length $L/\eta_{r}\sim\lambda_{s}$, while that of PBC corresponds to $g=0$.
@@
where $0<x_{1,2}<L$. $B$ and $B^{\prime}$-terms are the reflected waves of $A$ and $A^{\prime}$-terms respectively. After the reflection, the real parts of momenta switch, but their imaginary parts change due to the SOC. As a result, the imaginary parts of $k_{1,\downarrow}$ and $k_{2,\uparrow}$ are at the order of $1/L$,
@@
while that of the reflected momenta $k_{1,\uparrow}$ and $k_{2,\downarrow}$ become finite as
@@
Here $\chi_{i}=2\pi n_{i}$ with $i=1,2$. $A^{\prime}$ and $B^{\prime}$-terms are the transmitted waves of $A$ and $B$-terms respectively. The PBC yields $A^{\prime}/A=e^{ik_{1,\downarrow}L}$ and $B^{\prime}/B=e^{ik_{2,\downarrow}L}$.
We view $A$, $A^{\prime}$-terms as the incident waves and $B$, $B^{\prime}$-terms as the reflected waves. Since the imaginary parts of their momenta behave differently, in the case of $L\gg\lambda_{s}$, the reflected waves can be dropped if the inter-particle distance exceeds $\lambda_{s}$. Then the wavefunction is simplified to the product of plane-waves and a Jastrow factor,
@@
where $e^{-\frac{1}{2}W(x_{1}-x_{2})}$ is given by the sum of step-functions modified by the imaginary parts of the complex momenta
@@
$W(r)=2\eta_{r}\left(\frac{r}{L}+\theta(-r)\right),\ \ -L<r<L.$ (11)
The exact and approximated wavefunctions Eq. (7) and Eq. (10) are shown in Fig. 1 respectively. The spin-up particle is fixed at $x_{2}\!=\!0$. Increasing $x_{1}$ from 0, $\varphi(x_{1},x_{2})$ rapidly reaches the peak located at $x_{1,\rm peak}\sim\ln(L\lambda_{s}^{-1})/(L\lambda_{s}^{-1})$. If $L\gg\lambda_{s}$, $x_{1,\rm peak}$ coincides with $x_{2}$. The peak is followed by a slow decay at the length scale of $L$, which means the spin-down particle prefers the right side of the spin-up one. This can be understood as a weaker version of the NHSE with localization length $L/\eta_{r}$. Note that this length decreases as $g$ increases, indicating that the localization is enhanced by the repulsive interaction, which distinguishes it from the conventional NHSE suppressed by repulsions as reported in the previous studies [29, 27, 33, 12, 47]. The behavior at $x_{1}<0$ can be obtained by applying the PBC. As we will explain later, the two-body wavefunction can be generalized to the many-body case where the above picture still holds.
Figure 1: The 2-body wavefunction $\varphi(x_{1},x_{2})$ with particles of opposite spins by fixing the spin-up one at $x_{2}=0$. The parameter values are $\chi_{1}=\chi_{2}=0$, $m=1$, $\alpha=1$ and $L=50$. The peak appears at $x_{1}\sim\ln(L\lambda_{s}^{-1})/(L\lambda_{s}^{-1})$. In the case of $L\gg\lambda_{s}$, the peak is located at $x_{1}=x_{2}$ and $\varphi$ becomes discontinuous. It implies that the spin-down particle tends to lie on the right side of the spin-up particle, forming a resonant pair on the ring. The wavefunction becomes more localized as the strength of repulsive interaction $g$ increases, which indicates that the resonance is enhanced by the repulsive interaction.
Many-body problem: BA equations.— For eigenstates with $N_{\downarrow}$ down spins and $N_{\uparrow}=N-N_{\downarrow}$ up spins, the corresponding BA equations are:
@@
Note that the scattering is always non-diffractive in terms of $\overline{k}_{l}$, which ensures the integrability of the model.
The solutions to BA equations are significantly simplified due to the non-Hermitian SOC. For example, in the two-body case, the solutions can be effectively expressed by the Eq. (8), since those reflected waves with momenta $\{k_{1,\uparrow},k_{2,\downarrow}\}$ are removed in the approximation to the wavefunctions. The same thing happens in the many-body case. As illustrated in Fig. (2), the many-body wavefunction is composed of many ‘plane waves’, connected with each other by scattering. Those ‘plane waves’ in the same column are connected by transmission, which share the same momenta since transmission exchanges the positions of the particles while preserving their momenta. For example, $A_{2}$, $A_{3}\cdots$ are the transmission descendants of $A_{1}$, with momenta
$k_{i,\downarrow}$ $=$ $(\chi_{i}+i\mathcal{\eta}_{r}N_{\uparrow})/L,~~k_{j,\uparrow}=(\chi_{j}-i\mathcal{\eta}_{r}N_{\downarrow})/L.$ (14)
Here $1\leq i\leq N_{\downarrow}$, $N_{\downarrow}+1\leq j\leq N$, $\chi_{l}=2\pi n_{l}$ and $\mathcal{\eta}_{r}$ is defined in Eq. (6). $A_{1}$, $A_{2}$, $A_{3}\cdots$ are denoted as incident waves, like the $A$ and $A^{\prime}$-terms in the two-body case. Each incident wave has many reflected descendants, aligned in the same row. Due to the diffractive reflection described by Eq. (9), these descendants decay much faster than the incident waves, therefore can be discarded in the dilute limit defined as,
@@
where $d=L/N$ is the average inter-particle distance. Detailed calculations are found in the SM. II. Note that our approximated solution exhibits a singularity at $\alpha=0$, since the dilute limit is broken in that case.
@@
Figure 2: The construction of the Bethe-ansatz wavefunction. It is composed of many ‘plane waves’. Each plane wave is a product of Slater determinants, due to the Fermi statistics. The ‘plane waves’ in the same column or the same row are connected by transmission or reflection respectively. Due to the diffractive reflection described by Eq. (9), the reflected descendants of $A_{i}$ are all suppressed.
Figure 3: The Monte-Carlo simulations of the average ‘spin dipole’ $\langle p\rangle$ at $L=100$ with different particle numbers. As $\eta_{r}=\ln(1+g/\alpha)$ increases, $\langle p\rangle/L$ jumps from zero to a finite value, indicating that the transition is of 1st order. ($a$) and ($b$) shows the configuration of the two phases. If $\alpha$ is fixed, the more localized ($b$) phase emerges only when $g$ is big enough, which implies that the many-body resonance is enhanced by the repulsive interaction.
@@
$\!\!\!\varphi(x_{\downarrow};x_{\uparrow})\!=\!\text{det}(e^{\frac{i\chi_{i_{1}}x_{i_{2}}}{L}})\text{det}(e^{\frac{i\chi_{j_{1}}x_{j_{2}}}{L}})e^{-\frac{1}{2}\sum_{ij}\!W(x_{ij})},$ (16)
in which two $\text{det}(\cdots)$ represent the Slater determinants of the plane-waves for spin-up and down particles respectively; the two-body Jastrow factor in Eq. (10) is also generalized to the many-body case. Here $1\leq i,i_{1},i_{2}\leq N_{\downarrow}$ and $N_{\downarrow}+1\leq j,j_{1},j_{2}\leq N$ are the coordinate indexes for spin-down and up particles respectively, and $x_{ij}=x_{i}-x_{j}$ represents the distance between them. This wavefunction is similar to that of the Hubbard model at $U\rightarrow\infty$ [30], where the BA wavefunction is factorized into a Slater determinaint of spinless fermions and a BA wavefunction of the spin-1/2 chain.
The above simplification is justified as follows. As illustrated in Fig. 2, the wavefunction is approximated to a summation of incident waves with different coordinate permutations. In other words,
$\varphi(x_{\downarrow};x_{\uparrow})$ $=\text{det}(e^{\frac{i\chi_{i_{1}}x_{i_{2}}}{L}})\text{det}(e^{\frac{i\chi_{j_{1}}x_{j_{2}}}{L}})e^{-\sum_{ij}\frac{\eta_{r}x_{ij}}{L}}$
$\times\sum_{Q}\theta\left(x_{Q_{1}}>x_{Q_{2}}>\cdots>x_{Q_{N}}\right)A(Q),$
where $Q$ represents the permutation $x_{Q_{1}}>x_{Q_{2}}>\cdots>x_{Q_{N}}$ with $1\leq Q_{l}\leq N$. $\sum_{Q}$ denotes the sum over all permutations. These incident waves differ by a ‘phase shift’, whose module is not $1$ due to the non-Hermitian SOC. After switching a pair of neighboring particles with opposite spins, the amplitudes are changed by
@@
which is momentum independent at the leading order. As proved in SM. II, in this case the summation of step functions can be organized into
@@
$\frac{N-2}{4},-\frac{N-6}{4},\cdots-1,0,1,\cdots\frac{N-6}{4},\frac{N-2}{4}.$
@@
$\mathcal{H}=\sum_{i_{1}<i_{2}}V(x_{i_{1},i_{2}})+\sum_{j_{1}<j_{2}}V(x_{j_{1},j_{2}})+\sum_{ij}W(x_{ij}),$ (19)
where $x_{i_{1},i_{2}}=x_{i_{1}}-x_{i_{2}}$ and $x_{j_{1},j_{2}}=x_{j_{1}}-x_{j_{2}}$. $V(r)=-2\ln\left|\sin\frac{\pi}{L}r\right|$ originates from the Pauli exclusion principle, which describes an effective repulsion between particles of identical spins; $W$ brings an unidirectional attraction between opposite spins, with spin-up particles preferring the left side of spin-down ones.
If $V$ dominates, particles tend to uniformly distribute along the ring, with a weak pairing tendency between opposite spins to take the advantage of $W$, as depicted in Fig. 3 ($a$). Conversely, if $W$ dominates, the system prefers phase separation, such that nearly all spin-up particles lie on the left of spin-down particles, making the configuration in Fig. 3 ($b$) more favorable. The competition between $V$ and $W$ is investigated via Monte-Carlo simulations with the probability distribution Eq. (18). We consider the “dipole” strength $p=\frac{1}{(N/2)^{2}}\sum_{ij}p_{ij}$, where $p_{ij}$ is the dipole between a spin-down and up particle located at $x_{i}$ and $x_{j}$ respectively,
$p_{ij}=\left\{\begin{array}[]{l}x_{ij},~~~~~~~~~~~~~~~~~|x_{ij}|\leq L/2,\\ x_{ij}-\operatorname{sgn}(x_{ij})L,\ |x_{ij}|>L/2.\end{array}\right$. (20)
With this definition, $L/2<p\leq L/2$. The thermodynamics average $\langle p\rangle$ is plotted in Fig. 3. As $\eta_{r}$ increases, $\langle p\rangle/L$ evolves from zero, which is consistent with Fig. 3 ($a$), to a finite value, illustrated in Fig. 3 ($b$). The transition takes places at $\eta_{r}\approx 3.15$. Note that $\eta_{r}$ measures the strength of resonance, which is enhanced by the repulsive interaction as shown in the two-body case. Therefore, the emergence of the ($b$) phase demonstrates a many-body ‘skin effect’ driven by the repulsive interaction, which is in stark contrast to the conventional NHSE [29, 27, 33, 12, 47].